OpenStudy (anonymous):

Select the true statement about the graph of f(x) = -2x^3 + 6X^2 + 18X -18.

OpenStudy (anonymous):

I'm assuming that the first two can be eliminated because it's saying that the minimum is postive, while the maximum is negative? Is that possible?

OpenStudy (anonymous):

You know that from the sign of the largest term that the cubic is negative at large positive x and positive at large negative x so the minimum should be to the left of the maximum

OpenStudy (anonymous):

min (-1,-28) max(3,36)

OpenStudy (anonymous):

Oooh. So I was waaay off. Lol.

OpenStudy (anonymous):

No your logic was sound. Your minima were larger than the maxima. On the graph it cannot be. Good instincts. The rest you might have to plug and check the last two answers.

OpenStudy (anonymous):

Oh, okay Lol. But, how would I plug in to check the answers?

OpenStudy (anonymous):

surdawi, what's the graph of?

OpenStudy (anonymous):

your equation

OpenStudy (anonymous):

Ah, I see it now

OpenStudy (anonymous):

f(-1) = -2(-1)^3 + 6(-1)^2 + 18(-1) -18 f(-1) = +2 + 6 -18 -18 f(-1) = +8 -36 = -28

OpenStudy (anonymous):

Ahhh, okay. Thank you both for helping me! (: